OpenStudy (anonymous):

What is the vertex of the parabola given by y = x2 − 10x + 17?

OpenStudy (anonymous):

-b/2a 10/2

OpenStudy (anonymous):

a=1(from the x^2), b=-10, c=17

OpenStudy (anonymous):

y= x^2 - 10x + 17 complete the square to find (h,k) y= (x^2 -10x + _ ) - _ +17 y= (x^2 -10x + 25) - 25 + 17 factor out what's in the parentheses y= (x-5)^2 - 8 and from here you can get your vertex (h,k) (5,-8)

OpenStudy (anonymous):

thankyou! haha i was confused :P

OpenStudy (anonymous):

Your welcome :)